OpenStudy (unklerhaukus):

Solve the Differential Equation: \[y^\prime\left(2x+y^2\right)=y\]

OpenStudy (unklerhaukus):

\[y^\prime\left(2x+y^2\right)=y\]\[\left(2x+y^2\right)\cdot\text dy=y\cdot\text dx\]\[\frac{\text dx}{\text dy}=\frac {2x}y+y\]\[x^\prime-\frac 2y x=y\]

OpenStudy (unklerhaukus):

\[R(y)=e^{\int{-\frac{2}y}\text dy}=e^{-2\ln|y|}=y^{-2}=\frac 1{y^2}\] \[\left(\frac x{y^2}\right)^\prime=\frac{y}{y^2}\] \[\frac x{y^2}=\int \frac 1y\text dy\] \[\frac {x}{y^2}=\ln y^2+c_1\] \[x=y^2\ln y+c_1y^2\] \[y^2\ln y -x=cy^2\]

OpenStudy (unklerhaukus):

there is a mistake on the third line from the bottom \(\ln y^2\)**\(\ln y\)

OpenStudy (rogue):

Wow, awesome! :D never saw a differential equation being solved this way.

OpenStudy (unklerhaukus):

is there another way to solve this question?

OpenStudy (rogue):

There might be, but I just recently learned basic DE's, so I don't really know of another method :(